OpenStudy (anonymous):

Assume that w=f(s^3+t^2) and f'(x)=e^x. Find partial w/partial t and partial w/partial s

OpenStudy (amistre64):

since f'(x) = e^x ; f(x) = e^x as well

OpenStudy (amistre64):

w = e^(s^3+t^2)

OpenStudy (anonymous):

how did that happen?

OpenStudy (amistre64):

w=f(x) w' = f'(x) x' you know what f' is, so integrate it to determine f

OpenStudy (anonymous):

and that will give me the partial derivative i seek

OpenStudy (amistre64):

that will give you the explicit form of w in terms of s and t plug is s^2 + t^2 for x ; then derive

OpenStudy (anonymous):

okay let me look over this for a bit, thanks

OpenStudy (amistre64):

w = f(s^3+t^2) = e^(s^3+t^2)